# Causal Regime Detection in Energy Markets With Augmented Time Series Structural Causal Models

### v1

@@ 1 Introduction
Energy markets exhibit intricate causal relationships between weather conditions, generation technologies, grid constraints, and price formation. Unlike traditional financial markets, electricity cannot be stored economically, creating immediate causal dependencies between supply, demand, and pricing ziel2018day. Current electricity price forecasting methods focus on prediction accuracy without providing causal interpretation or counterfactual reasoning capabilities lago2021forecasting.
Recent advances in counterfactual reasoning for high-dimensional data through Augmented Structural Causal Models (ASCM) enable principled causal inference in complex domains pan2024counterfactual; pan2025counterfactual. However, these frameworks assume static causal structures and cannot handle the temporal dependencies and regime changes inherent in energy systems.
@@ 2 Related Work
Causal Inference in Financial Markets. Causal data science has been applied to financial stress testing and risk management gao2018causal; rigana2024navigating. However, energy markets present unique challenges due to physical constraints, storage limitations, and renewable intermittency that require specialized treatment.
Electricity Price Forecasting. Traditional approaches focus on statistical methods (ARIMA, VAR) or machine learning models (LSTM, gradient boosting) for price prediction lago2021forecasting; ziel2018day. While achieving good predictive performance, these methods lack explicit causal interpretation and cannot answer counterfactual queries.
Counterfactual Reasoning. ASCM frameworks pan2024counterfactual; pan2025counterfactual enable counterfactual reasoning in high-dimensional spaces but assume static causal structures. Time series causal methods like TNCM-VAE thumm2025towards, SCIGAN bica2020estimatingthe and Causal Transformer melnychuk2022causal address temporal dynamics but do not handle regime changes in learned causal graphs.
@@ 3.1 Energy Market Causal Structure
Let $\mathcal{G}_{t}$ represent the time-varying directed acyclic graph (DAG) encoding causal relationships at time $t$. A causal regime change occurs when:
@@
where ${\mathcal{M}}(t)$ represents the causal mechanisms (functional relationships and noise distributions) governing the DAG at time $t$, also known as Structural Causal Model (SCM) pearl2009causality.
@@ 3.2 ATSCM for Energy Markets
We extend ASCM pan2024counterfactual; pan2025counterfactual to temporal energy data through a three-level generative hierarchy with learned causal structure:

### v2

@@ 1 Introduction
Energy markets exhibit intricate causal relationships between weather conditions, generation technologies, grid constraints, and price formation. Unlike traditional financial markets, electricity cannot be stored economically, creating immediate causal dependencies between supply, demand, and pricing Ziel and Weron (2018). Current electricity price forecasting methods focus on prediction accuracy without providing causal interpretation or counterfactual reasoning capabilities Lago et al. (2021).
Recent advances in counterfactual reasoning for high-dimensional data through Augmented Structural Causal Models (ASCM) enable principled causal inference in complex domains Pan and Bareinboim (2024, 2025). However, these frameworks assume static causal structures and cannot handle the temporal dependencies and regime changes inherent in energy systems.
@@ 2 Related Work
Causal Inference in Financial Markets. Causal data science has been applied to financial stress testing and risk management Gao et al. (2018); Rigana et al. (2024). However, energy markets present unique challenges due to physical constraints, storage limitations, and renewable intermittency that require specialized treatment.
Electricity Price Forecasting. Traditional approaches focus on statistical methods (ARIMA, VAR) or machine learning models (LSTM, gradient boosting) for price prediction Lago et al. (2021); Ziel and Weron (2018). While achieving good predictive performance, these methods lack explicit causal interpretation and cannot answer counterfactual queries.
Counterfactual Reasoning. ASCM frameworks Pan and Bareinboim (2024, 2025) enable counterfactual reasoning in high-dimensional spaces but assume static causal structures. Time series causal methods like TNCM-VAE Thumm and Mijares (2025), SCIGAN Bica et al. (2020) and Causal Transformer Melnychuk et al. (2022) address temporal dynamics but do not handle regime changes in learned causal graphs.
@@ 3.1 Energy Market Causal Structure
Let $\mathcal{G}_{t}$ represent the time-varying directed acyclic graph (DAG) encoding causal relationships at time $t$. A causal regime change Huang et al. (2024) occurs when:
@@
where ${\mathcal{M}}(t)$ represents the causal mechanisms (functional relationships and noise distributions) governing the DAG at time $t$, also known as Structural Causal Model (SCM) Pearl (2009).
@@ 3.2 ATSCM for Energy Markets
We extend ASCM Pan and Bareinboim (2024, 2025) to temporal energy data through a three-level generative hierarchy with learned causal structure: